OpenStudy (anonymous):

all solution lie between -10 and 10 x^4-3x^3+7x-2=0 the solution set is ?

OpenStudy (anonymous):

HINT:... x^4-3x^3+7x-2=0 x^4-3x^3 + 9/4 x^2 = 9/4 x^2 -7x +2 (x^2 - 3/2 x)^2 + 2*k*(x^2 - 3/2 x) /2 +k^2/4 = 2*k*(x^2 - 3/2 x) /2 +k^2/4 -7x +2 +9/4 x^2 (x^2-3/2 x +k/2)^2 =(9/4 + k)x^2 - (7+3k/2)x +k^2/4 + 2 the value of k is so chosen such that RHS becomes a perfect square... and hence solve the tow reulting quadratic...

OpenStudy (anonymous):

Thanks